Consider the following ecnomy:

Total Population = 1000

Total Adult Civilian (non-institutionalized) Population = 700

Number of people employed = 200

Number of people without jobs but looking for work = 80

What is the labor force participation rate in this economy?

Question 11 options:

28%

40%

70%

none of the above
Answer #1

Labor force participation rate = Labor force / Adult population * 100

Labor force = employed + unemployed = 200 + 80 = 280

Labor force participation rate = 280 / 700 * 100 = 40%